OpenStudy (anonymous):

Prove (cotØ+1)/ cscØ = cosØ + sinØ

OpenStudy (solomonzelman):

can you work on both sides, or work only with one side ?

OpenStudy (anonymous):

only one

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf \frac{cotØ+1}{cscØ}=cosØ+sinØ }\) being that cscØ=1/sinØ therefore 1/cscØ = 1/(1/sinØ)=sinØ So... \(\LARGE\color{blue}{ \bf (cotØ+1)sinØ=cosØ+sinØ }\) so far so good ?

OpenStudy (anonymous):

yes it is good

OpenStudy (solomonzelman):

So right now we have, \(\LARGE\color{blue}{ \bf (cotØ+1)sinØ=cosØ+sinØ }\) lets expand the left side, \(\LARGE\color{blue}{ \bf cotØ~sinØ~+sinØ=cosØ+sinØ }\) Now, cotØ =cos Ø/ sinØ So, substituting cos Ø/ sinØ instead of cotØ \(\LARGE\color{blue}{ \bf \frac{cosØ }{sinØ} ~sinØ~+sinØ=cosØ+sinØ }\) now the sines in the fraction cancel, \(\LARGE\color{blue}{ \bf \frac{cosØ }{\cancel{ sinØ} } ~\cancel{ sinØ} ~+sinØ=cosØ+sinØ }\) need more help ?

OpenStudy (anonymous):

wow that looks like 10 times better than what i was trying to do, thanks so much!!!

OpenStudy (solomonzelman):

Anytime, BTW you are yellow now :)

OpenStudy (anonymous):

Aww thanks!!